OpenStudy (anonymous):

find m....3/4m - 1/3 =1/2

OpenStudy (anonymous):

3/4m = 5/6 4m = 18/5 m =9/10

OpenStudy (anonymous):

thank u

OpenStudy (gw2011):

Do you mean (3/4)m-1/3=1/2? If so, then the solution is as follows: (3/4)m=(1/2)+(1/3) (3/4)m=5/6 m=(5/6)(4/3) m=20/18=10/9